# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v2

@@ I Introduction
Recent advances in deep learning (DL) have brought about a tectonic shift in the seismic interpretation workflow [1, 2], mirroring broader trends across other domains like geoscience [3, 4], sustainable energy systems [5, 6, 7] and biomedical applications [8, 9, 10, 11, 12, 13, 14]. DL-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [15, 16]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [17, 18]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific neural networks, feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [19, 20]. In general, the growing availability of DL methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [19, 20].
@@
Two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  1, orange box) and the strategies used to adapt models across them (Fig. 1, green box). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [21], and eventually provide weak performance for the interpretation task.
@@
To address the challenge of domain adaptation, researchers have proposed a variety of methods that address the effects of domain differences along the pipeline. Several studies introduce enhancements to model architectures [22], dataset construction [23, 24, 15, 25], uncertainty estimation [26] or explore training strategies aimed at improving data efficiency and transferability [27, 28]. A common approach involves pretraining models on synthetic data and subsequently fine-tuning them on real seismic volumes, as in [15]. More recent methods have explored using self-supervised learning to learn robust features without large amounts of labeled data [29, 30, 31, 32, 33, 34], or adapting large vision foundation models trained on natural images to the seismic domain [20, 35, 36]. While these approaches offer promising results, their effectiveness is often dependent on the compatibility between source and target domains, and they remain vulnerable to issues such as catastrophic forgetting [37] during adaptation. Moreover, the success of these approaches is influenced by several factors such as architecture and training design choices, as well as dataset-specific properties such as seismic section resolution, heterogeneity, and sample size.
@@
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 1 (purple box). Fault labels are derived from expert interpretation and are often subjective [38, 39], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [40, 19]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
This paper presents a large-scale benchmarking study regarding the performance variability of finetuning strategies for fault delineation models across training regimes, data characteristics, and architectural choices. These fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [41], using randomly initialized models, jointly training on multiple fault datasets, and using standard domain adaptation (DA) strategies. Our contributions are as follows:
  • We present the first large-scale benchmarking study focused on fault delineation under domain shift, spanning over 200 experimental setups across 8 DL architectures, 5 pretraining-finetuning datasets and 3 evaluation ones, as well as several training configurations.
@@
The remainder of this paper is organized as follows. Section II reviews prior work on machine learning for fault delineation in the context of seismic interpretation workflows. Section III describes the three datasets used in our benchmarking study, highlighting their geological and statistical characteristics. Section IV details our experimental setup, including data preparation, model architectures, and evaluation protocols. Section V presents the results of over 200 training configurations, analyzed across generalization and transferability, training dynamics, and metric evaluation. Section VI introduces an additional analysis on geometric and topological metrics to characterize dataset–model interactions. Finally, Section VII concludes the paper by summarizing key findings, outlining practical guidelines, and identifying open challenges for future research.
@@ II-A Data Preparation and Preprocessing
In computer vision segmentation pipelines, raw images from natural scenes [42, 43, 44] and medical scans [45, 46, 47] are first standardized to ensure consistent spatial dimensions and intensity distributions [48, 49]. Typical preprocessing steps include resizing or cropping to a fixed height and width, data augmentation, whitening [50], and contrast adjustments to mitigate variability in lighting or sensor settings [51, 52]. These operations guarantee that each input conforms to the network’s architectural requirements and that learned features are not biased by extraneous intensity fluctuations.
Seismic segmentation extends these practices to volumetric data acquired in formats such as SEG-Y (.sgy), raw binary (.dat), or serialized NumPy arrays (.npy). A common workflow begins by reading trace headers to assemble a 3D volume of dimensions (inline $\times$ xline $\times$ time/depth), and either processing the volume at that level [15, 19] or extracting 2D inline or crossline sections for further processing [23, 32]. To enlarge the training set and fit GPU memory constraints, each section is tiled with an overlapping sliding window of size $H\times W$ pixels (the effects of different tiling standards are studied and results presented in Section V-B1.).
@@ II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [45], DeepLab [43], and SegFormer [53] typically accept fixed‐size $n\times n$ patches (usually $n\in\{128,256,512\}$), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. The binary cross-entropy (BCE) loss [54] focuses on classifying each pixel correctly, while the Dice loss [55] measures the overlap between predicted and ground-truth fault regions, making it well-suited for imbalanced datasets. In practice, many works combine BCE and Dice to take advantage of both pixel fidelity and shape alignment [56, 57].
In seismic fault delineation, similar principles apply but with additional considerations. Input patches are typically larger to capture fault continuity across sections, and training often uses overlapping windows with a stride smaller than the patch size to ensure boundary faults are adequately sampled [15, 58, 59, 60]. The aforementioned loss functions, coupled with learning rate schedules and different regularization functions, form the backbone of seismic training setups [61, 62, 63]. By carefully tuning window sizes and loss compositions, researchers mitigate class imbalance and preserve structural continuity in fault delineation [64].
@@ II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [65, 66, 67]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [68, 69, 70, 71, 72]. Domain adaptation encompasses strategies to reduce this gap when the target data is limited, for instance by aligning feature distributions between source and target domains [73, 74]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [75, 76].
@@
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is especially true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [77, 61]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [15]) often fail to transfer effectively to real data from different basins. Recent surveys [78, 79] provide comprehensive overviews over the way deep learning approaches in exploration seismology struggle with such domain shifts and emphasize fault segmentation as a particularly brittle task.
@@
To mitigate these challenges, researchers have explored a variety of domain adaptation strategies, including feature-space [80] and frequency-space [81] alignment, adversarial learning [82], and seismic-style transfer [83]. Fault-specific adaptations include synthetic fault injection to enrich training distributions [84], contrastive learning schemes designed for curvilinear structures [34], and self-supervised pretraining frameworks that improve transferability across surveys with limited labels [19, 20]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [85, 86] or large-scale seismic simulations with domain-specific decoders [61, 15]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [37], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.
@@ II-D Inferential Behavior
In semantic segmentation, inference entails mapping learned feature representations to discrete pixel‐level predictions [42]. The fidelity of this mapping depends on the architecture’s ability to aggregate context and fuse features: models with narrow receptive fields may accurately localize edges but miss global structure [45], while those with extensive context capture coherent regions at the expense of sharp boundaries [43]. During inference, the bias–variance trade‐off emerges as a tension between boundary precision and region consistency, sometimes mitigated by post‐processing, e.g., Conditional Random Fields [87] or edge‐aware refinement modules [88].
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [77]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [45], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [61]. DeepLab’s [43] atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries enhancing continuity across sections, though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [53].
Furthermore, recent works in both seismic [78, 79] and general DL [89, 90] domains emphasize that inferential biases are not merely architectural, but also dataset-driven: synthetic datasets may thus encourage smoother, well-connected predictions, while field datasets can induce models to replicate jagged, discontinuous, or crossover-heavy structures (we analyze these trends in depth using the fault characteristic metrics described in Section II-F). New approaches attempt to counteract these biases by introducing geological priors, like fault continuity preservation [20] or curvature-based losses [34]. These studies highlight that inferential behavior in seismic fault delineation is shaped jointly by architectural design and the structural biases present in training data, underscoring the importance of evaluation frameworks that go beyond pixel overlap to account for geological plausibility.

### v3

@@ I Introduction
Recent advances in deep learning (DL) have brought about a tectonic shift in the seismic interpretation workflow [isah_review_2025, wenxue_overview_2025], mirroring broader trends across other domains like geoscience [yaqoob_geocrack_2024, qaiser_vision_2025], sustainable energy systems [yaqoob_advancing_2025, yaqoob_microcrystalnet_2025, yaqoob_fluidnet-lite_2025] and biomedical applications [logan2022multi, logan2022patient, kokilepersaud2023clinically, quesada2022mtneuro, kokilepersaud2022gradient, iffa2025development, ansari_survey_2025]. DL-assisted approaches are leveraged in different parts of the pipeline, specifically in automated fault detection. Faults are geologically critical features that control fluid flow in Earth’s subsurface, influence reservoir compartmentalization, and pose drilling hazards in hydrocarbon exploration [wu2019faultSeg, chowdhury2025unified]. Beyond the energy sector, faults play a central role in earthquake nucleation and propagation, as well as geohazard and risk assessment in tectonically active regions [benzion2008collective, KANAMORI20031205]. Accurate and scalable fault interpretation is therefore a high-impact task across several geoscience domains. Early work in this area often repurposed semantic segmentation models from computer vision (such as those developed for natural images) to detect discontinuities in seismic sections. However, these generic frameworks struggled with the unique characteristics of seismic data, particularly the thin, curvilinear geometry of faults and the presence of acquisition artifacts and structural noise. As a result, the field has increasingly shifted toward fault delineation, a task-specific variant of segmentation that emphasizes the extraction of coherent fault structures rather than general pixel-wise classification. This shift has prompted the development of domain-specific neural networks, feature encoders, adapted loss functions, and structural priors that better capture the morphological signatures of faults [rs16050922, faultsam]. In general, the growing availability of DL methods has accelerated the adoption of data-driven interpretations that have provided new insights into fault structures within large seismic volumes [rs16050922, faultsam].
@@
Two central and intertwined challenges in the interpretation pipeline arise from the diversity of data sources (Fig.  1, orange box) and the strategies used to adapt models across them (Fig. 1, green box). Seismic datasets differ widely in their geological characteristics, acquisition parameters, and resolution, among other factors. Synthetic data, while providing a fully specified ground truth, often fails to capture the complexity and variability of field data. Prior datasets from other regions, though more realistic, may still differ significantly from new target volumes. These domain shifts, whether from synthetic-to-real or across real-world basins, can cause pretrained models to generalize poorly to new sites [nasim2020seismic], and eventually provide weak performance for the interpretation task.
@@
To address the challenge of domain adaptation, researchers have proposed a variety of methods that address the effects of domain differences along the pipeline. Several studies introduce enhancements to model architectures [LI2023105412], dataset construction [alaudah2018structure, alaudah2019machine, wu2019faultSeg, AN2021107219], uncertainty estimation [benkert2022reliable] or explore training strategies aimed at improving data efficiency and transferability [benkert2023samples, mustafa2021man]. A common approach involves pretraining models on synthetic data and subsequently fine-tuning them on real seismic volumes, as in [wu2019faultSeg]. More recent methods have explored using self-supervised learning to learn robust features without large amounts of labeled data [wang2016interactive, cohen2006detection, roberts2001curvature, di2017seismic, shafiq2018towards, kokilepersaud2022volumetric], or adapting large vision foundation models trained on natural images to the seismic domain [faultsam, Quesada_2024_CVPR, quesada2024benchmarking]. While these approaches offer promising results, their effectiveness is often dependent on the compatibility between source and target domains, and they remain vulnerable to issues such as catastrophic forgetting [kirkpatrick2017overcoming] during adaptation. Moreover, the success of these approaches is influenced by several factors such as architecture and training design choices, as well as dataset-specific properties such as seismic section resolution, heterogeneity, and sample size.
@@
The third critical complication in the pipeline arises in evaluation, as depicted in Fig. 1 (purple box). Fault labels are derived from expert interpretation and are often subjective [alcalde2017impact, guillon2020ground], particularly in complex or ambiguous structural settings. This subjectivity leads to inconsistencies in ground truth across datasets or among annotators, and complicates objective model evaluation. Furthermore, a given metric may penalize correct but unannotated predictions, while failing to reflect geological plausibility or structural consistency [sarajarvi2020robust, rs16050922]. To address this, our study provides an in-depth analysis of commonly used evaluation metrics (such as Dice coefficient, Chamfer and Hausdorff distances) in the context of fault interpretation. This analysis highlights the need for a more holistic evaluation approach that integrates quantitative measures with geological interpretability.
This paper presents a large-scale benchmarking study regarding the performance variability of finetuning strategies for fault delineation models across training regimes, data characteristics, and architectural choices. Rather than introducing a new model, this work establishes a systematic experimental framework: a methodological foundation for reproducible and cross-domain evaluation of deep learning architectures in seismic fault interpretation. Our benchmarked fault delineation models include a combination of eight deep learning architectures, trained and fine-tuned on three distinct datasets, including both synthetic and real volumes. Our experiments cover a broad spectrum of training strategies, including pretraining on a single fault dataset, pre-training on large-scale natural image dataset such as ImageNet [deng2009imagenet], using randomly initialized models and jointly training on multiple fault datasets. Beyond benchmarking, we also use targeted case studies (through domain adaptation and continual learning setups) to diagnose model behavior under catastrophic forgetting and domain shift, demonstrating how the benchmark can serve as a tool for both evaluation and insight.
While prior works have examined transfer learning and domain adaptation in geoscience [sun2016deep, tsai2018learning, trinidad2023seismic], these efforts have typically been limited in scope: focusing on single architectures or isolated dataset pairs. By contrast, our study provides the first large-scale, systematic benchmark that unifies evaluation across multiple model families, datasets, and training paradigms. This effort parallels the role of domain-shift benchmarks in fields like medical imaging [guan_domain_2022], where standardized comparisons have proven essential for understanding generalization and reproducibility across domains.
Our contributions are as follows:
  • We present the first large-scale benchmarking study focused on fault delineation under domain shift, spanning over 200 exhaustive combinations of experimental setups across 8 DL architectures, 5 pretraining-finetuning datasets and 3 evaluation ones, as well as several training configurations.
@@
The remainder of this paper is organized as follows. Section II reviews prior work on machine learning for fault delineation in the context of seismic interpretation workflows. Section III describes the three datasets used in our benchmarking study, highlighting their geological and statistical characteristics. Section IV details our experimental setup, including data preparation, model architectures, and evaluation protocols. Section V presents the results of over 200 training configurations, analyzed across generalization and transferability, training dynamics, and metric evaluation. Section VI introduces an additional analysis on geometric and topological metrics to characterize dataset–model interactions. Finally, Section VII concludes the paper by summarizing key findings, outlining practical guidelines, and identifying open challenges for future research. The full cosde and data used for this paper can be found at https://github.com/olivesgatech/large-bench-geo.
@@ II-A Data Preparation and Preprocessing
In computer vision segmentation pipelines, raw images from natural scenes [long2015fully, chen2017deeplab, kokilepersaud2023focal] and medical scans [ronneberger2015u, isensee2021nnu, prabhushankar2022olives] are first standardized to ensure consistent spatial dimensions and intensity distributions [ioffe2015batch, ulyanov2016instance]. Typical preprocessing steps include resizing or cropping to a fixed height and width, data augmentation, whitening [prabhushankar2017generating], and contrast adjustments to mitigate variability in lighting or sensor settings [shorten2019survey, wang2017effectiveness]. These operations guarantee that each input conforms to the network’s architectural requirements and that learned features are not biased by extraneous intensity fluctuations.
Seismic segmentation extends these practices to volumetric data acquired in formats such as SEG-Y (.sgy), raw binary (.dat), or serialized NumPy arrays (.npy). A common workflow begins by reading trace headers to assemble a 3D volume of dimensions (inline $\times$ xline $\times$ time/depth), and either processing the volume at that level [wu2019faultSeg, rs16050922] or extracting 2D inline or crossline sections for further processing [alaudah2018structure, di2017seismic]. To enlarge the training set and fit GPU memory constraints, each section is tiled with an overlapping sliding window of size $H\times W$ pixels (the effects of different tiling standards are studied and results presented in Section V-B1.).
@@ II-B Training Dynamics and Setups
Popular network choices designed for natural image segmentation are sensitive to training and design choices like model architecture, input window size, and loss functions. Networks such as U-Net [ronneberger2015u], DeepLab [chen2017deeplab], and SegFormer [xie2021segformer] typically accept fixed‐size $n\times n$ patches (usually $n\in\{128,256,512\}$), balancing the need for sufficient context against GPU memory limits. During each training iteration, these patches are sampled, sometimes at multiple scales to expose the network to varied object sizes and to regularize against overfitting. Segmentation models optimize losses designed to reconcile pixel‐wise accuracy with region‐level coherence. The binary cross-entropy (BCE) loss [cox1958regression] focuses on classifying each pixel correctly, while the Dice loss [milletari2016v] measures the overlap between predicted and ground-truth fault regions, making it well-suited for imbalanced datasets. In practice, many works combine BCE and Dice to take advantage of both pixel fidelity and shape alignment [wazir2022histoseg, zu2024resaceunet].
In seismic fault delineation, similar principles apply but with additional considerations. Input patches are typically larger to capture fault continuity across sections, and training often uses overlapping windows with a stride smaller than the patch size to ensure boundary faults are adequately sampled [wu2019faultSeg, di2019developing, di2019improving, di2019semi]. The aforementioned loss functions, coupled with learning rate schedules and different regularization functions, form the backbone of seismic training setups [dou2022md, benkert2024effective, mustafa2023active]. By carefully tuning window sizes and loss compositions, researchers mitigate class imbalance and preserve structural continuity in fault delineation [mustafa2024visual].
@@ II-C Generalization and Transferability
Generalization refers to a model’s ability to maintain performance when presented with new, unseen data that (ideally) follows the same distribution as the training set [goodfellow2016deep, benkert2022example, mustafa2021joint]. In segmentation, this means accurately delineating objects or regions under variations in lighting, scale, or background texture. However, a common obstacle to generalization is domain shift, which occurs when the statistical properties of training and test data differ (such as changes in camera sensors or scene composition) often leading to degraded performance [torralba2011unbiased, Temel2017_NIPS, temel2018cure, temel2019traffic, prabhushankar2022introspective]. Domain adaptation encompasses strategies to reduce this gap when the target data is limited, for instance by aligning feature distributions between source and target domains [ganin2016domain, chen2019temporal]. Another common technique to adapt to target domains is fine-tuning: a network pretrained on a large, generic dataset (the source) is adapted to a more specialized task (the target) by retraining some or all layers, thereby leveraging learned representations while adjusting to new data characteristics [panigrahi2020survey, yosinski2014transferable].
@@
In seismic segmentation, generalization to new domains is particularly difficult due to variability in acquisition parameters, stratigraphy, frequency content, and noise levels across surveys. This is especially true for fault delineation, where subtle and discontinuous features are easily obscured by processing artifacts or geologic heterogeneity [ercoli2023evidencing, dou2022md]. Models trained on one domain (such as synthetic datasets like FaultSeg3D [wu2019faultSeg]) often fail to transfer effectively to real data from different basins. Recent surveys [mousavi_deep-learning_2022, mousavi_applications_2024] provide comprehensive overviews over the way deep learning approaches in exploration seismology struggle with such domain shifts and emphasize fault segmentation as a particularly brittle task.
@@
To mitigate these challenges, researchers have explored a variety of domain adaptation strategies, including feature-space [sun2016deep] and frequency-space [trinidad2023seismic] alignment, adversarial learning [tsai2018learning], and seismic-style transfer [du2022disentangling]. Fault-specific adaptations include synthetic fault injection to enrich training distributions [wang2022structural], contrastive learning schemes designed for curvilinear structures [kokilepersaud2022volumetric], and self-supervised pretraining frameworks that improve transferability across surveys with limited labels [rs16050922, faultsam]. Transfer learning pipelines typically rely on pretrained encoders (either from natural images [he2016deep, simonyan2014very] or large-scale seismic simulations with domain-specific decoders [dou2022md, wu2019faultSeg]) followed by fine-tuning on limited labeled sections. These strategies must navigate a tradeoff between plasticity and stability: aggressive weight updates are prone to inducing catastrophic forgetting [kirkpatrick2017overcoming], while conservative tuning may fail to capture domain-specific features. As a result, generalization in seismic fault delineation requires careful calibration of both model architecture and adaptation strategy to handle the subtle, spatially sparse structures across diverse geological, acquisition, and imaging settings.
@@ II-D Inferential Behavior
In semantic segmentation, inference entails mapping learned feature representations to discrete pixel‐level predictions [long2015fully]. The fidelity of this mapping depends on the architecture’s ability to aggregate context and fuse features: models with narrow receptive fields may accurately localize edges but miss global structure [ronneberger2015u], while those with extensive context capture coherent regions at the expense of sharp boundaries [chen2017deeplab]. During inference, the bias–variance trade‐off emerges as a tension between boundary precision and region consistency, sometimes mitigated by post‐processing, e.g., Conditional Random Fields [krahenbuhl2011efficient] or edge‐aware refinement modules [pohlen2017full].
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [ercoli2023evidencing]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [ronneberger2015u], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [dou2022md]. DeepLab’s [chen2017deeplab] atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries enhancing continuity across sections, though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [xie2021segformer].
Furthermore, recent works in both seismic [mousavi_deep-learning_2022, mousavi_applications_2024] and general DL [recht2019imagenet, geirhos2020shortcut] domains emphasize that inferential biases are not merely architectural, but also dataset-driven: synthetic datasets may thus encourage smoother, well-connected predictions, while field datasets can induce models to replicate jagged, discontinuous, or crossover-heavy structures (we analyze these trends in depth using the fault characteristic metrics described in Section II-F). New approaches attempt to counteract these biases by introducing geological priors, like fault continuity preservation [faultsam] or curvature-based losses [kokilepersaud2022volumetric]. These studies highlight that inferential behavior in seismic fault delineation is shaped jointly by architectural design and the structural biases present in training data, underscoring the importance of evaluation frameworks that go beyond pixel overlap to account for geological plausibility.